16:00: calibration of Hexapod to Patrol Cameras on the SX system: REC= 20160313_220000. This will be used for the fast laser guide w/ HVC green channel...

Truth sensing: when a new KL is available we must edit the REC list in the truesense.conf file located in /home/aoacct/AO/ and /home/flao/TEST/calib/wfs/current/truesense actually these file exist for the AOeng user but do not exist for the flao user...

SX BCU: trimmed to 3.5V

SX PnCCD: pumped again...

20:00: BCU re-installed on SX, failing PCD installed on DX.

We try to collimate the telescope in binocular mode.

NGC2419

Preset sent as mag 14, clouds are really thick!

Hip40075

We exercise the paddle GUI, Steve sends us a brief summary of its usage.

Hip58175

23:12: preset as mag 10.2, since there are still too many clouds we close the AO loop just on FLAO BIN4

  • Combo REC, FLAO only (BIN4): 20160313_063600

TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20160314_063929 1.5,0.5, 0 334.1 20151212_233600 -1 F F
20160314_064112 1.5,0.6, 0 323.178 20151212_233600 -1 F F
20160314_064256 2,0.8, 0 322.649 20151212_233600 -1 F F
20160314_064416 4,1.5, 0 339.392 20151212_233600 -1 F F
20160314_064538 5,1.5, 0 329.853 20151212_233600 -1 F F
Then we crash and we must restart the loop, we're guiding on the LGS:

TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20160314_070044 2,0.8, 0 335.56 20151212_233600 -1 F F
20160314_070133 3,1.2, 0 495.999 20151212_233600 -1 F F
20160314_070205 3,1.2, 0 710.152 20151212_233600 -1 F F
20160314_070619 3,1.2, 0 368.185 20151212_233600 -1 F F
20160314_071543 3,1.5,0.3 32.0061 20160313_063600 -1 T F
20160314_071628 3,1.5,0.3 29.6797 20160313_063600 -1 T F
We try the fast guiding w/ green channel:

TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20160314_072917 3,1.5,0.3 35.7334 20160313_063600 -1 T F
20160314_073021 3,1.5,0.3 35.7628 20160313_063600 -1 T F
We also build a FLAO only REC BIN2:

  • Combo REC, FLAO only (BIN2): 20160313_075000

1:00: general SW restart, AARB, AOS, AdSec are failing...

We continue to test the fast guiding

Hip71389

2:30: we acquire cubes of patrol cameras images w/ and w/o VibComp, files are located into /tmp/GOX/patCam/:

  • 20160314_120118 (w/o)
  • 20160314_120122 (w/)

20160314_patrol_camera_vibration_compensation_on.jpg

???

5:30: new target, preset as 13.8, we acquire new patCams cubes, much less wind:

  • 20160314_124646

TN AO Gain Res. WF [nm rms] AO rec DIMM ["] JCL PCL Notes
20160314_125230 0.2,0.1,0.05 546.937 20160312_032600 -1 T F
20160314_125303 0.5,0.35,0.1 627.49 20160312_032600 -1 T F

-- MarcoBonaglia - 22 Mar 2016
Topic revision: r1 - 22 Mar 2016, MarcoBonaglia
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback